WARNING: This web is not used anymore. Please use PDBService.MinuteS24November05 instead!
 

Minutes 24 November 05

  • VRVS/phone: Stefan (ATLAS), Gordon (RAL), Sasha(ATLAS), Yuri Smirnov(BNL), Vladimir Kravchuck(RIPN), Marc Rodriguez(PIC), Pierre-Etienne(IN2P3), Marek Sarvas, Andrea Capriotti(CNAF)
  • @CERN: Radovan(POOL), Birger(ARDA), Dirk (3D), Eva (3D), Saima(CMS)

  • Sites and hardware setup for '06 production (Dirk)
    • Target date: March 2006
    • Proposed 2006 setup schedule (see transparencies from agenda, slide 14)
    • Proposed Tier1 Service setup (see transparencies from agenda, slide 15)
    • BNL: single instances for ATLAS setup, investigating FibreChannel storage setup for cluster
    • IN2P3: Solaris Cluster, 2 nodes (1 node per experiment), type V480, storage is not a problem, licence discussion afterwards
    • PIC: setting up databases schedule to January, possibly join production phase in March: Dirk will discuss with Gonzalo offline
    • RAL: 2 x 2 nodes clusters, no major problem with schedule
    • Dirk will follow the discussion offline with the sites not attending the meeting today. GridKA, CNAF were connected but could not speek (or hear?)
  • 3D testbed status:
    • Replication was stuck at the beginning of the week
    • Problem: lot of spilled LCRs (changes not consumed) on disk due to one site is down since long time, no memory available to capture new changes
    • Solution: open a TAR, Oracle recommendation is to remove the site down from the testbed
    • Replication is working now, but it will be necessary some time to be up to date with the changes
  • COOL test status
    • Stefan was able to install COOL application after some installation trouble maybe due to hard coded installation path - Dirk suggests to Stefan to report the bug to COOL/SPI for further investigation
    • Tests against CERN, Oxford and RAL run successfully
    • Limited network is the bottleneck right now
    • no test of replication yet- will start in 1 week
    • proposed introduce variations at some time: increase the workload to see which will be the limitation
  • CMS online-offline replication test status
    • Streams setup is ready
    • Saima is looking at the materialized views replication - DDL changes are ignored by Streams replication
  • AMGA replication test status
    • Streams setup for one schema between the 2 databases at CERN on the 3D testbed (Eva and Luis)
    • Birger started to fill the AMGA schema on the source database at CERN with 1 million entries spread over 100 tables
    • Only 4% of the entries had been replicated when the replication got stuck (because of the problem described in the section testbed status above)
    • Replication is happening now after solving the problem, Birger should check again if all the entries have been replicated
  • General recommendation is to send notification of performance test plans to the 3d project lists to be able to plan and schedule the tests in order to avoid disturbing between tests
  • Dynamic Deployment of Database Services presentantion by Sasha (see transparencies from agenda)

Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r3 - 2005-11-29 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback